# Fractional diffusion as the limit of a short range potential Rayleigh gas

### v1

@@ 5.1. Estimate the gain term of the collision operator
Let $f\in L^{1}(\mathds{R}^{3})$ a radial and unimodal distribution with $f(v)(1+|v|)\in L^{1}(\mathds{R}^{3})$, then
$\int_{\mathds{R}^{3}}f(v_{1}^{\prime})|v_{1}-v_{1}^{\prime}|^{\lambda}\,\mathrm{d}v_{1}^{\prime}=\int_{\mathds{R}^{3}}f(v_{1}^{\prime})|v_{1}|^{\lambda}\,\mathrm{d}v_{1}^{\prime}+o(|v_{1}|^{\lambda}),$ (5.2)
for large $|v_{1}|$, where $\lambda:=\frac{n-5}{n-1}<0$ for $n\in(2,5)$.
By taking now the full integral $K(f)(v_{1})$ we have
$K(f)(v_{1})$
$=\int_{\mathds{R}^{3}}\left[\frac{2}{|v_{1}-v_{1}^{\prime}|^{2}}\int_{E_{v_{1},v_{1}^{\prime}}}g_{0}(v_{2}^{\prime})B(2v_{1}-v_{1}^{\prime}-v_{2}^{\prime},\omega)\,\mathrm{d}v_{2}^{\prime}\right]f(v_{1}^{\prime})\,\mathrm{d}v_{1}^{\prime}$
$=4\pi C_{2}\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}|v_{1}|^{-q}\int_{\mathds{R}^{3}}f(v_{1}^{\prime})\bigg{[}C_{S}|v_{1}-v_{1}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{1}^{\prime}|^{\frac{n-7}{n-1}})\bigg{]}\,\mathrm{d}v_{1}^{\prime}$
$=4\pi C_{2}\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}|v_{1}|^{-q}\bigg{[}C_{S}\int_{\mathds{R}^{3}}f(v_{1}^{\prime})|v_{1}-v_{1}^{\prime}|^{\frac{n-5}{n-1}}\,\mathrm{d}v_{1}^{\prime}+\int_{\mathds{R}^{3}}f(v_{1}^{\prime})\mathcal{O}(|v_{1}-v_{1}^{\prime}|^{\frac{n-7}{n-1}})\,\mathrm{d}v_{1}^{\prime}\bigg{]}$
$=\underbrace{4\pi C_{2}\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}C_{S}}_{=:C_{K,a}}\bigg{[}|v_{1}|^{-q+\frac{n-5}{n-1}}+o(|v_{1}|^{-q+\frac{n-5}{n-1}})\bigg{]},$
for large $|v_{1}|$. Here we used Proposition 5.1 and the fact that $f$ is a probability distribution. Therefore, when $\Theta\in(0,\frac{\pi}{2}-\sigma)$ the operator $K(f)$ is behaving as
$K(f)(v_{1})$ $=C_{K,a}\bigg{[}|v_{1}|^{-q+\frac{n-5}{n-1}}+o(|v_{1}|^{-q+\frac{n-5}{n-1}})\bigg{]},$ (5.3)
for $|v_{1}|$ large, $q\in(4,\frac{4n}{n-1})$ and $n\in(2,5]$.
Proof of Proposition 5.1.
@@
We first note that $f\in L^{\infty}(\mathds{R}^{3}\setminus B_{1}(0))$ where $B_{1}(0)$ is the unit ball as $f$ is a radial function which decays monotonically to 00.
For $\lambda:=\frac{n-5}{n-1}$, $n\in(2,5)$, that is $-3<\lambda<0$, consider first the integral on the ball $B(v_{1},\frac{|v_{1}|}{2})$, i.e.
@@
which is finite since $f\in L^{1}(\mathds{R}^{3})$. We used the fact that the term $\frac{1}{|{\rm e}_{1}-\frac{v_{1}^{\prime}}{|v_{1}|}|^{|\lambda|}}$ is bounded by $2$ on the set $\Omega\setminus B(v_{1},\frac{|v_{1}|}{2})$. This completes the proof of the proposition. ∎
5.2. Estimate the loss term of the collision operator
@@
$\nu(v_{1})$ $=\int_{\mathbb{S}^{2}}\int_{\mathds{R}^{3}}g_{0}(v_{1}^{\prime})B(v_{1}-v_{1}^{\prime},\omega)\,\mathrm{d}v_{1}^{\prime}\,\mathrm{d}\omega$
@@
$=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\int_{\mathbb{S}^{2}}\int_{\mathds{R}^{3}}g_{0}(v_{1}^{\prime})\bigg{[}|v_{1}-v_{1}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{1}^{\prime}|^{\frac{n-7}{n-1}})\bigg{]}\Theta(\sin\Theta)^{-1}\,\mathrm{d}v_{1}^{\prime}\,\mathrm{d}\omega$
$=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\int_{\mathbb{S}^{1}}\int_{\mathds{R}^{3}}g_{0}(v_{1}^{\prime})\bigg{[}|v_{1}-v_{1}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{1}^{\prime}|^{\frac{n-7}{n-1}})\bigg{]}\int_{0}^{\tilde{\sigma}}\Theta\,\mathrm{d}\Theta\,\mathrm{d}v_{1}^{\prime}\,\mathrm{d}\psi$
$=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\pi\tilde{\sigma}^{2}\int_{\mathds{R}^{3}}g_{0}(v_{1}^{\prime})\bigg{[}|v_{1}-v_{1}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{1}^{\prime}|^{\frac{n-7}{n-1}})\bigg{]}\,\mathrm{d}v_{1}^{\prime}$
@@
$=\underbrace{\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\pi\tilde{\sigma}^{2}}_{=:\tilde{C}_{K}}\bigg{[}\left[|v_{1}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}|^{\frac{n-7}{n-1}})\right]\underbrace{\int_{\mathds{R}^{3}}g_{0}(v_{1}^{\prime})\,\mathrm{d}v_{1}^{\prime}}_{=1}+o(|v_{1}|^{\frac{n-5}{n-1}})\bigg{]}.$

### v2

@@ 5.1. Estimate of the loss term of the collision operator
Let $f\in L^{1}(\mathds{R}^{3})$ be a radial and unimodal distribution. Suppose in addition that $f(v)(1+|v|)\in L^{1}(\mathds{R}^{3})$. Then, for any $\lambda<0,$
$\int_{\mathds{R}^{3}}f(v_{1}^{\prime})|v_{1}-v_{1}^{\prime}|^{\lambda}\,\mathrm{d}v_{1}^{\prime}=\int_{\mathds{R}^{3}}f(v_{1}^{\prime})|v_{1}|^{\lambda}\,\mathrm{d}v_{1}^{\prime}+o(|v_{1}|^{\lambda}),\quad\text{as}\ |v_{1}|\to\infty.$ (5.5)
Proof.
@@
We first note that $f\in L^{\infty}(\mathds{R}^{3}\setminus B_{1}(0))$ where $B_{1}(0)$ is the unit ball as $f$ is a radial function which decays monotonically to 0.
Consider first the integral on the ball $B(v_{1},\frac{|v_{1}|}{2})$, i.e.
@@
which is finite since $f\in L^{1}(\mathds{R}^{3})$. We used the fact that the term $\frac{1}{|{\rm e}_{1}-\frac{v_{1}^{\prime}}{|v_{1}|}|^{|\lambda|}}$ is bounded by $2$ on the set $\Omega\setminus B(v_{1},\frac{|v_{1}|}{2})$. This completes the proof of the lemma. ∎
Proposition 5.2.
The function $\nu$ is bounded and is behaving as
$\nu(v_{1})$ $=\tilde{C}\bigg[|v_{1}|^{\frac{n-5}{n-1}}+o(|v_{1}|^{\frac{n-5}{n-1}})\bigg],$ (5.7)
as $|v_{1}|\to\infty$, $q\in(4,\frac{4n}{n-1})$ and $n\in(3,5]$.
Proof.
@@
[…]
then boundedness for bounded $v_{1}$ follows from the boundedness of $\mathcal{B}$ implied by Lemma 3.2 and Corollary 3.4. The estimates follow from
[…]
@@
$=\frac{2^{\frac{2}{n-1}}}{a^{2}}\int_{\mathbb{S}^{1}}\int_{\mathds{R}^{3}}g_{0}(v_{1}^{\prime})\bigg[|v_{1}-v_{1}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{1}^{\prime}|^{\frac{n-7}{n-1}})\bigg]\int_{0}^{\frac{\pi}{2}}\Theta\,\mathrm{d}\Theta\,\mathrm{d}v_{1}^{\prime}\,\mathrm{d}\psi$
$=\frac{2^{\frac{2}{n-1}}}{a^{2}}\pi\frac{\pi}{2}^{2}\int_{\mathds{R}^{3}}g_{0}(v_{1}^{\prime})\bigg[|v_{1}-v_{1}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{1}^{\prime}|^{\frac{n-7}{n-1}})\bigg]\,\mathrm{d}v_{1}^{\prime}$
@@
$=\underbrace{\frac{2^{\frac{2}{n-1}}}{a^{2}}\pi\frac{\pi}{2}^{2}}_{=:\tilde{C}}\bigg[\left[|v_{1}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}|^{\frac{n-7}{n-1}})\right]\underbrace{\int_{\mathds{R}^{3}}g_{0}(v_{1}^{\prime})\,\mathrm{d}v_{1}^{\prime}}_{=1}+o(|v_{1}|^{\frac{n-5}{n-1}})\bigg].$